OpenStudy (anonymous):

Explain what happens if you don't get a square root completely simplified on the first try. How is this like simplifying fractions?

OpenStudy (anonymous):

You can just keep repeating the process. For example, \sqrt{100} = \sqrt{4*25} = \sqrt{2*2*25} = 2\sqrt{25}However, that isn't completely simplified, as 5*5=25, so we can continue:\sqrt{100} = 2\sqrt{25} = 2\sqrt{5*5} = 2*5 = 10 10 is the fully simplified form, as the prime factorization of 10 is 2*5, and none of the factors are repeated.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np